One number is greater than thrice the other number by 2. If 4 times the smaller
number exceeds the greater by 5, find the numbers

Answer by stanbon(57246)

One number is greater than thrice the other number by 2. If 4 times the smaller
number exceeds the greater by 5, find the numbers
-----------------------
Equation:
x = 3*y + 2
4y = x + 5
----
3y + 2 = 4y - 5
y = 7
--
x = 3y+2 = 23
------------
Cheers,
Stan H.